A mountain climbing expedition establishes a base camp and two intermediate camps, A and B. Camp A is 13,500 m east of and 3,000 m above base camp. Camp B is 8000 m east of and 1800 m higher than Camp A. Determine the displacement between base camp and Camp B
1
Expert's answer
2021-09-16T10:26:02-0400

Find the total vertical and horizontal displacement:


"H=13500+8000=21500\text{ m},\\
V=3000+1800=4800\text{ m},\\
D=\sqrt{V^2+H^2}=22030\text{ m}."